• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 18:23:28 on March 2,2006

    Entry number 164828

    keyword=Optical fiber situation.

    Ameya and I checked and cleaned all the long optical fibers.
    There are two fibers no signal at all (F5 and F8). Four good ones
    (F1,F2,F3 and F7) now connected to the four lasers (single 1, 2 and DUO 1, 2). The remain two spare fibers, one (F6) has signal with laser point, will need to check with laser before using it (need a short laser controlled access). The other one (F4, SN009) has a mark
    on the upstair side. Will talk to the company to see if it is possible to clean without damaging it.
    Of the 5-to-1 fiber, one has a burn mark in the middle of the metal jacket, but its tranmission is fine so far. Should watch and check
    often.